A ball bounces 2/3 of the distance through which it falls. If the second rebound is 72 cm, the height,in cm,through which the ball originally dropped was
(A)32 (B)48 (C)108 (D)162 (E) None of these

Answer by stanbon(75887)

A ball bounces 2/3 of the distance through which it falls. If the second rebound is 72 cm, the height,in cm,through which the ball originally dropped was
-----------------
Let the original fall distance be "x":
Equation:
[(2/3)^2]x = 72
(4/9)x = 72
(1/9)x = 18
x = 162 cm
---------------------
Cheers,
Stan H.
---------------------




(A)32 (B)48 (C)108 (D)162 (E) None of these